OpenStudy (anonymous):

Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y = one halfx y = one halfx + 6 y = one halfx + 9

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@ganeshie8 @jim_thompson5910

OpenStudy (anonymous):

@Mertsj @texaschic101

OpenStudy (anonymous):

I don't understand your question.

OpenStudy (anonymous):

please help im in a hurry

OpenStudy (mertsj):

1. Distribute the 1/2 on the right 2. Add 3 to both sides.

OpenStudy (anonymous):

@Mertsj 1/2x + 6 1/2

OpenStudy (mertsj):

What is: \[\frac{1}{2}(x+6)\]

OpenStudy (anonymous):

if you distribute the 1/2 you get 1/2x together ?

OpenStudy (mertsj):

\[\frac{1}{2}(x+6)=\frac{1}{2}x+3\]

OpenStudy (anonymous):

ok and then

OpenStudy (anonymous):

@Mertsj ?

OpenStudy (mertsj):

1. Distribute the 1/2 on the right 2. Add 3 to both sides.

OpenStudy (mertsj):

Do step 2

OpenStudy (anonymous):

3 1/2x+6

OpenStudy (mertsj):

Do you know that "add 3" means +3?